1. On December 31, year 0, ABC purchased a machine in exchange for an interest-bearing note requiring 7 payments of $287457 at the end of each year. The first payment was made on December 31, year 1. At the date of the transaction, the prevailing rate of interest for this type of note was 5.2%. The initial value of the machine is:

Answer: ____________

2. ABC purchased a machine that has an estimated useful life of 6 years for $4834

  • The machine's salvage value is estimated to be $468
  • ABC uses the straight-line depreciation method

The depreciation for the second year of the machine's life is: __________
Answer #1

1. Initial value of the machine = Present value of annual payments

= 287457*5.7445

= $1651440

2. Depreciation for the second year = (4834-468)/6 = $427.67 or $428